OpenStudy (anonymous):

Find the thirty-second term of the following sequence. 9, 15, 21, ...

OpenStudy (anonymous):

Answer Is 195

OpenStudy (anonymous):

a+(n-1)d

OpenStudy (priyar):

wait..first do u know A.P formula ?

OpenStudy (priyar):

it is : \[a _{n}=a +(n-1)d\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

195 is correct

OpenStudy (priyar):

where \[a _{n} \] is the nth term.. n is the no. of terms..a is the first term..d is the difference between first two terms

OpenStudy (priyar):

here a=9 d=(15-9)=6 n=32.. substitute these values in the above eq to find \[a_{32}\] thats it..!

OpenStudy (priyar):

u got it @Bodrul_Ahmed ?